OpenStudy (anonymous):

Kevin has a collection of nickels, dimes and quarters whose total is $6.50. There are twice as many dimes as nickels and 8 less quarters than nickels. How many of each coin does Kevin have? Please Show Your Work! Thank You!!

OpenStudy (anonymous):

@Hero can you help me plz!!

hero (hero):

.05n + .10d + .25q = 6.50 d = 2n q = n - 8

OpenStudy (anonymous):

Thx!!

hero (hero):

That's not the answer. That's just a hint.

OpenStudy (anonymous):

i know.

OpenStudy (anonymous):

@Hero i think i'm doing it wrong. i got 4.14

hero (hero):

.05n + .10(2n) + .25(n-8) = 6.50 .05n + .20n + .25n - 2 = 6.50 (.05 + .20 + .25)n = 6.50 + 2 .50n = 8.50 n = 8.50/.50

hero (hero):

Do you get it bro?

OpenStudy (anonymous):

i forgot to multiply 0.25 * 8 i'm so stupid! Thx for helping!